# Machine Learning to Predict Spectral Anisotropy in Valence-to-Core X-ray Emission Spectroscopy

### v1

@@ 2.4 Additional Input Features
This section focuses on defining each additional descriptor beyond the anisotropy-matrix–derived quantities and motivating its inclusion in the model. As we have shown, the QAMS alone is not a strong predictor of the SAMS, which motivates incorporating additional local geometric and chemical factors. Along this line, we include the density, the normed space group number, the number of 3dd-electrons, standard deviation and average electronegativity of the cluster as additional factors. The correlation between all factors, along with the SAMS, is displayed in Figure 4.
@@
The normalized space group number ($N_{SG}$), as shown in Eq. 9,
@@
where $N_{s}$ is the space group number of the structure, and $N_{t}=230$ is the total number of crystallographic space groups. Eq. 9 provides a rough, simple analogue for the symmetry of the crystal system and, to a lesser extent, the symmetry of the local environment around the metal.
The number of $3d$ electrons on the central atom is given in Eq. 10,
@@
where $Z_{c}$ is the atomic number of the central atom and $q_{c}$ is its charge, as determined by the oxidation state guesser. The constant $18$ accounts for the electrons in the filled inner shells ($1s^{2}\,2s^{2}\,2p^{6}\,3s^{2}\,3p^{6}$), and this parameter is intended to add additional chemical information in a way that is analogous to oxidation state.
The average and standard deviation of electronegativity were added as features to capture the distribution of atomic bonding tendencies within each cluster. These quantities were computed using the Pauling electronegativity values provided by the Pymatgen periodic table [Ong.etal2013]. For each cluster, we evaluated the electronegativity of all constituent atoms and then calculated the mean and standard deviation, which together characterize both the central tendency and variability of electronegativity in the system. The density of the structure is obtained from the Materials Project [Jain2013].
@@ 2.5 Dataset curation
Crystallographic Information Files (CIFs) were queried from the Materials Project in their primitive cell format [Jain2013]. We selected structures that are experimentally observed and that contain at least one 3dd transition metal atom, resulting in a dataset of approximately 11,500 compounds. Figure 6 (a) shows the distribution of transition metals within the dataset, illustrating that certain elements, such as Cu and Fe, are overrepresented relative to others like Sc or Ti, reflecting trends in the Materials Project database. Figures 6 (b) and (c) further show the distributions of the QAMS and SAMS metrics respectively. While not identical for every element, the QAMS and SAMS distributions for the same transition metal are qualitatively similar. The most spectrally isotropic element is the Sc, with very few compounds at all that exhibit any significant anisotropy. As we will see in Section 3, the limitations of the Sc subset (both the limited number of compounds and the narrow distribution) will make it difficult to for the model to sufficiently generalize. When splitting the dataset into training and testing subsets, this elemental imbalance is mitigated by selecting a proportionally representative number of structures for each transition metal.
@@ 2.6 Corvus Workflow and FEFF
Corvus is an automation workflow handler [Kas.etal2021a] which uses a CIF and user defined information to write a FEFF input file. FEFF is a real-space Green’s function code that employs a full multiple-scattering formalism [FEFF9.6] for calculating the polarized VtC-XES. The following procedure is laid out in a diagram in Figure 5. Corvus starts by using pymatgen [Ong.etal2013] to predict the oxidation states of all the atoms in a unit cell, which are in turn used to construct the weighted quadrupole matrix as defined in Eq. 3a. An important note is that for crystal structures that contain multiple instances of the same 3dd transition metal, the quadrupole tensors are averaged over all sites before diagonalizing. The orthogonal basis that comes from diagonalizing the quadrupole is then used as the basis vectors for the subsequent FEFF VtC-XES calculation, and the generation of all subsequent features.
FEFF input files are programmatically generated with individual clusters around each transition metal element in the unit cell. The SCF and FMS radii are dynamically set to contain the closest 30 atoms to the central atom in each calculation, which helps achieve consistent convergence. All calculations were performed in the presence of a core hole, consistent with previous theoretical treatments of VtC-XES and the final-state rule [mortensen-2017]. It should be noted that we also limit the number of unique potentials so that the absorbing atom in the center of the cluster is assigned one unique potential, while all other atoms have a potential which is only unique to their atomic number. The polarized output spectra are used to calculate the SAMS according to Eq. 1 and 2 to produce a target property that our model tries to predict. The total computation time across the full dataset took approximately 3300 hours of CPU time.
@@
Download CIFs from Materials Project Estimate oxidation states of atoms using Pymatgen Build out clusters, diagonalize quadrupole tensor. Formulate the QAMS, use the eigenvectors as an orthogonal basis for the spectral calculation. FEFF Polarized VtC-XES Calculate integrated spectral differences according to Eq.1 and Eq.2. Compile the SAMS together into a complete dataset
@@ 2.7 Machine Learning
In this work, we employ a random forest regressor (RFR) [breiman2001random] to predict the SAMS of 3dd transition metal clusters. An RFR is an ensemble learning method in which many decision trees are trained on bootstrapped samples of the data [glassner-2021], and the final prediction is obtained as the average of their outputs. Each tree is built through recursive binary splitting of the feature space, where only a randomly selected subset of features is considered at each split, thereby reducing correlation among trees and improving model robustness [james2023isl, hastie2009elements, loh2014fifty]. This model architecture provides a practical compromise between accuracy, interpretability, and computational cost, making it well suited for relating the constructed features to the anisotropy observed in VtC-XES. The RFR implementation was taken from scikit-learn [sklearn-rfr].
2.7.1 Model Training and Hyperparameter Optimization
@@ 3 Results & Discussion
3.0.1 Model Results
We use the coefficient of determination $R^{2}$ [chicco2021rsq] to evaluate model performance along with the mean absolute error (MAE). The $R^{2}$ metric quantifies the fraction of variance in the target variable explained by the model predictions, while the MAE provides an alternative performance metric through the average absolute prediction error.
@@ 3.0.1 Model Results
Our model achieves a high predictive accuracy, and the small difference between training and testing performance, $\Delta R^{2}=-0.0421$ and $\Delta\mathrm{MAE}=+0.0020$, indicates that the model generalizes well to unseen data. The relatively high $R^{2}$ score ($0.81$ on the testing set) demonstrates that the model captures the majority of the variance in the SAMS from the chosen chemical and geometric input features. This level of predictive performance is consistent with other recent applications of machine learning in materials spectroscopy and property prediction, where $R^{2}$ values in the range of 0.70.70.90.9 are typically reported for regression tasks involving complex, correlated physical quantities (e.g. chemical interpretation of near-edge x-ray absorption structure) [torrisi-2020, iwayama-2022, guda-2021].
@@
Figure 7 shows the results of the trained model run on the test dataset. The vertical axis is the predicted SAMS value while the horizontal axis is the true SAMS value. We see that the distribution is strongly skewed along the diagonal, which is consistent with the high $R^{2}$ performance. Most of the materials are weakly anisotropic or not anisotropic at all, leading to a large density of points near the origin of the plot. Despite its strong overall performance, some systematic issues are evident. The model tends to slightly under predict SAMS values at the higher end of the distribution, while predictions for smaller SAMS values exhibit higher density and tighter clustering. This bias likely reflects the imbalance in the available data (Figure 6), where extreme high-anisotropy configurations are less common.
3.0.2 Model Interpretability
RFR feature contributions can be interpreted using variable importance measures [strobl-2007], and in this work we focus on evaluation via permutation feature importance (PFI). We chose this measure because impurity-based feature importance can be biased, especially for features with many distinct values [strobl-2007]. This is relevant for our data, where SAMS values can take any real number between 0 and 1. Impurity-based measures also depend on training-set statistics and therefore may not reflect predictive performance on unseen data.
PFI measures the importance of a feature by quantifying the change in model performance after the feature has been randomly permuted. In simple terms, this means shuffling the feature’s values among all data points to break its relationship with the target and observing how much the model worsens. If performance degrades substantially, the feature is considered more important. A more detailed discussion of PFI can be found in [molnar2025].
@@ 3.0.2 Model Interpretability
A feature ($j$) is considered important if the average importance, $\overline{\text{PFI}}_{j}$, is large, as this indicates a substantial degradation in performance when the feature is permuted [molnar2025interpretable]. However, the notion of “large” depends on the scale of the evaluation metric. Because the raw magnitudes of $\overline{\text{PFI}}_{j}$ depend on the dataset and model, they are not directly comparable across problems. To provide a meaningful scale, we normalize by the importance of a random noise feature, $\overline{\text{PFI}}_{\text{noise}}$, to provider a noise-normalized $\widetilde{\text{PFI}}_{j}$ in Eq. 11.
@@
A $\widetilde{\text{PFI}}_{j}\approx 1$ implies that feature $j$ is no more predictive than random noise, while $\widetilde{\text{PFI}}_{j}>1$ indicates predictive power beyond noise. For example, $\widetilde{\text{PFI}}_{j}=5$ means that feature $j$ is five times as informative as the noise baseline. As shown in Figure 8, the anisotropy-based descriptors IAMS and QAMS dominate the model’s predictive performance, consistent with their direct link to polarization-dependent spectral asymmetries discussed in Sec. 2. The modest importance of geometric and chemical descriptors such as the normalized space group, density, and DAMS indicates that the model captures secondary contributions from the structural environment. The slight advantage of IAMS over QAMS likely reflects reduced uncertainty in quantities derived directly from computed spectra compared to oxidation-state–dependent charge estimates. e
3.0.3 Model Performance for Each Individual Metal
Figure 9: Per-element performance of the random forest regressor evaluated on the test set. Ordered from largest (top) to smallest (bottom) atomic number. (a) shows the coefficient of determination $R^{2}$ computed separately for each metal using the combined (unbalanced) training dataset, while (b) shows the corresponding $R^{2}$ values obtained using the balanced dataset. In both panels, the model is trained on all elements and evaluated on element-specific subsets of the test data.
To better understand the limitations of the model, we analyzed its performance on a per-element basis. Specifically, we computed the coefficient of determination, $R^{2}$, for each individual metal using subsets of the test data corresponding to a single element, while keeping the random forest regressor (RFR) trained on the full dataset. The resulting per-element $R^{2}$ values are shown in Figure. 9 (a).
For clarity, we explicitly define the coefficient of determination as
@@ 3.0.3 Model Performance for Each Individual Metal
Among all elements, Sc exhibits the poorest performance, with an $R^{2}$ value close to zero, indicating that the model provides little improvement over a naive mean predictor for this element. Figure 10 shows predicted versus true SAMS values for the Sc (a) and Cu (b) subsets of the test data. Showcasing the difference between a subset of the data that the model performs well on and one that it does not.
For Sc, the coefficient of determination is $R^{2}\approx 0.01$, which corresponds to a ratio of squared errors $\approx 0.99.$ This indicates that the model reduces the error relative to a mean predictor by only about 1%. The small but positive $R^{2}$ value suggests that the model predictions are slightly shifted relative to the Sc-specific mean, consistent with a bias toward the global mean $\overline{y}_{\mathrm{comb,true}}$ of the full multi-element dataset.
In contrast to Sc, Cu (b) exhibits excellent performance, with an $R^{2}$ of $0.851$, demonstrating that the model can accurately learn patterns when sufficient representative data is present.
Next, we investigate if the poor performance of Sc may be due to under representation of this element in the materials project dataset. To address this, we constructed a balanced data set with an equal number of compounds (444 compounds from each element, see Figure 6 (a)). The training and testing sets were also split such that there was an even distribution of each element, with 354 samples in the training, and 90 samples in the test set. The results of this are presented in Figure. 9 (b), where the performance for Sc significantly worsens, suggesting that its poor predictive behavior likely stems from intrinsic differences in the characteristics of Sc compared to the other 3dd transition metals rather than data imbalance.
@@ 4 Conclusion
In this work we demonstrated that a random forest model is able to accurately predict a continuous anisotropy parameter using structural and chemical information extracted from crystallographic structures. This result provides an important step in creating new opportunities for machine learning driven materials discovery, as well as the design of materials with targeted spectroscopic properties. The model generalizes well to any crystalline 3dd transition metal structure (with the exception of Sc), but could easily be adapted to treat molecular systems as well by replacing or removing the normed space group feature.
@@
Future work includes extending our machine learning model to predict anisotropy in other x-ray methods, such as XAS. In particular, many of the structural and chemical descriptors that drive anisotropy in XES are also expected to influence absorption edge features, suggesting that the framework developed here could be transferable across spectroscopies. It may be possible to apply a transfer learning approach so that the already trained XES anisotropy model can be fine-tuned for predicting anisotropy in a different technique, which could dramatically reduce the computational cost compared to training a new model from scratch.

### v2

@@ 2.4 Additional Input Features
This section focuses on defining each additional descriptor beyond the anisotropy-matrix–derived quantities and motivating its inclusion in the model. As we have shown, the QAMS alone is not a strong predictor of the SAMS, which motivates incorporating additional local geometric and chemical factors. Along this line, we include the mass density, the normalized space group number, the number of 3dd-electrons, average and standard deviation of the electronegativity of the cluster as additional factors. The correlation between all factors, along with the SAMS, is displayed in Figure 4.
@@
The normalized space group number ($N_{SG}$), as shown in eq 9,
@@
where $N_{s}$ is the space group number of the structure, and $N_{t}=230$ is the total number of crystallographic space groups. Equation 9 provides a rough, simple analogue for the symmetry of the crystal system and, to a lesser extent, the symmetry of the local environment around the metal.
The number of $3d$ electrons on the central atom is given in eq 10,
@@
where $Z_{c}$ is the atomic number of the central atom and $q_{c}$ is its charge, as determined by the Pymatgen python library oxidation state guesser. The constant $18$ accounts for the electrons in the filled inner shells ($1s^{2}\,2s^{2}\,2p^{6}\,3s^{2}\,3p^{6}$), and this parameter is intended to add additional chemical information in a way that is analogous to oxidation state.
The average and standard deviation of electronegativity were added as features to capture the distribution of atomic bonding tendencies within each cluster. These quantities were computed using the Pauling electronegativity values provided by the Pymatgen periodic table Ong.etal2013. For each cluster, we evaluated the electronegativity of all constituent atoms and then calculated the mean and standard deviation. The mass density of the structure is obtained from the Materials Project Jain2013.
@@ 2.5 Dataset curation
Crystallographic Information Files (CIFs) were queried from the Materials Project in their primitive cell format Jain2013. We selected structures that are experimentally observed and that contain at least one 3dd transition metal atom, resulting in a dataset of approximately 11,500 compounds. Figure 6a shows the distribution of transition metals within the dataset, illustrating that certain elements such as Cu and Fe are overrepresented relative to others like Sc or Ti, reflecting trends in the Materials Project database. Figures 6b,c further show the distributions of the QAMS and SAMS metrics respectively. While not identical for every element, the QAMS and SAMS distributions for the same transition metal are qualitatively similar. The most spectrally isotropic element is Sc, with very few compounds exhibiting any significant anisotropy. As we will see in Section 3, the limitations of the Sc subset (both the limited number of compounds and the narrow distribution) make it difficult for the model to sufficiently generalize to all 3dd transition metals. When splitting the dataset into training and testing subsets, this elemental imbalance is mitigated by selecting a proportionally representative number of structures for each transition metal.
@@ 2.6 Corvus Workflow and FEFF
Corvus is an automation workflow handler Kas.etal2021a which uses a CIF and user defined information to write a FEFF input file. FEFF is a real-space Green’s function code that employs a full multiple-scattering formalism FEFF9.6 for calculating the polarized VtC-XES. The following procedure is laid out in a diagram in Figure 5. Corvus starts by using Pymatgen Ong.etal2013 to predict the oxidation states of all the atoms in a unit cell, which are in turn used to construct the weighted quadrupole matrix as defined in eq 3a. An important note is that for crystal structures that contain multiple instances of the same 3dd transition metal, the quadrupole tensors are averaged over all sites before diagonalizing. The orthogonal basis that comes from diagonalizing the quadrupole tensor is then used as the basis vectors for the subsequent FEFF VtC-XES calculation, and the generation of all subsequent features.
FEFF input files are programmatically generated for each individual cluster. The self-consistent field and full-multiple scattering radii used for the FEFF calculations are dynamically set to contain the closest 30 atoms to the central atom in each calculation, which helps achieve consistent convergence. All calculations were performed in the presence of a core hole, consistent with previous theoretical treatments of VtC-XES and the final-state rule mortensen-2017. It should be noted that we also limit the number of unique potentials so that the absorbing atom in the center of the cluster is assigned one unique potential, while all other atoms have a potential which is only unique to their atomic number. The polarized output spectra are used to calculate the SAMS according to eqs 1 and 2 to produce a target property that our model tries to predict. The total computation time across the full dataset took approximately 3300 hours of CPU time.
@@
Download CIFs from Materials Project Estimate oxidation states of atoms using Pymatgen Build out clusters, diagonalize quadrupole tensor. Formulate the QAMS, use the eigenvectors as an orthogonal basis for the spectral calculation. FEFF Polarized VtC-XES Calculate integrated spectral differences according to eq1 and eq2. Compile the SAMS together into a complete dataset
@@ 2.7 Machine Learning
In this work, we employ a random forest regressor (RFR) breiman2001random to predict the SAMS of 3dd transition metal clusters. An RFR is an ensemble learning method in which many decision trees are trained on bootstrapped samples of the data glassner-2021, and the final prediction is obtained as the average of their outputs. Each tree is built through recursive binary splitting of the feature space, where only a randomly selected subset of features is considered at each split, thereby reducing correlation among trees and improving model robustness james2023isl, hastie2009elements, loh2014fifty. This model architecture provides a practical compromise between accuracy, interpretability, and computational cost, making it well suited for relating the constructed features to the anisotropy observed in VtC-XES. The RFR implementation was constructed using the scikit-learn python library sklearn-rfr.
2.8 Model Training and Hyperparameter Optimization
@@ 3 Results & Discussion
3.1 Model Results
We use the coefficient of determination $R^{2}$ chicco2021rsq to evaluate model performance along with the mean absolute error (MAE). The $R^{2}$ metric quantifies the fraction of variance in the target variable explained by the model predictions, while the MAE provides an alternative performance metric through the average absolute prediction error.
@@ 3.1 Model Results
As shown in Table 3 our model achieves a high predictive accuracy, and the small difference between training and testing performance, $\Delta R^{2}=-0.0591$ and $\Delta\mathrm{MAE}=+0.0042$, indicates that the model generalizes well to unseen data. The relatively high $R^{2}$ score ($0.79$ on the testing set) demonstrates that the model captures the majority of the variance in the SAMS from the chosen chemical and geometric input features. This level of predictive performance is consistent with other recent applications of machine learning in materials spectroscopy and property prediction, where $R^{2}$ values in the range of 0.70.70.90.9 are typically reported for regression tasks involving complex, correlated physical quantities (e.g. chemical interpretation of near-edge x-ray absorption structure) torrisi-2020, iwayama-2022, guda-2021.
@@
Figure 7 shows the results of the trained model run on the test dataset. The vertical axis is the predicted SAMS value while the horizontal axis is the true SAMS value. We see that the distribution is strongly skewed along the diagonal, which is consistent with the high $R^{2}$ performance. Most of the materials are weakly anisotropic or not anisotropic at all, leading to a large density of points near the origin of the plot. Despite its strong overall performance, some systematic issues are evident. The model tends to slightly under predict SAMS values at the higher end of the distribution, while predictions for smaller SAMS values exhibit higher density and tighter clustering. This bias likely reflects the imbalance in the available data as shown in Figure 6, where extremely high-anisotropy configurations are less common.
3.2 Model Interpretability
RFR feature contributions can be assessed using variable importance measures strobl-2007. Here, we focus on permutation feature importance (PFI), which avoids biases present in impurity-based measures, particularly for features with many unique values. This is relevant for our data, where SAMS values can take any real number between 0 and 1. Impurity-based measures also depend on training-set statistics and therefore may not reflect predictive performance on unseen data. PFI quantifies a feature’s importance by measuring the change in model performance after randomly permuting its values. A more detailed discussion of PFI can be found in molnar2025. At a high level, PFI works by shuffling a feature’s values among all data points to break its relationship with the target and observing how much the model worsens. If performance degrades substantially, the feature is considered more important. Here the metric used to quantify the PFI is the negative mean squared error (NMSE), defined in Eq. 11,
$\text{NMSE}(y,\hat{y})=-\frac{1}{n}\sum_{i=1}^{n}\big(y_{i}-\hat{y}_{i}\big)^{2}$ (11)
where $y_{i}$ are the true values, $\hat{y}_{i}$ are the predicted values, and $n$ is the number of samples. We adopt NMSE following the convention in scikit-learn, where higher scores indicate better performance. Since the mean squared error is minimized, its negation ensures that values closer to zero represent better performance, while more negative values indicate worse performance sklearn_permutation_importance.
Let $\text{NMSE}_{0}$ denote a baseline score of the model on the dataset. For each feature $j$, we construct a permuted dataset $X^{\pi(j)}$ by randomly shuffling the entries in column $j$ while leaving all other columns unchanged. We then recompute the score as shown in Eq. 12.
$\text{NMSE}_{j}=\text{NMSE}(y,\hat{y}^{\pi(j)})$ (12)
The PFI for feature $j$ in a single permutation trial is defined in Eq. 13.
$\text{PFI}_{j}=\text{NMSE}_{0}-\text{NMSE}_{j}$ (13)
This process is repeated $R$ times with different random permutations, and the average importance is taken as in Eq. 14.
$\overline{\text{PFI}}_{j}=\frac{1}{R}\sum_{r=1}^{R}\Big(\text{NMSE}_{0}-\text{NMSE}_{j}^{(r)}\Big)$ (14)
@@ 3.2 Model Interpretability
A feature ($j$) is considered important if the average importance, $\overline{\text{PFI}}_{j}$, is large, as this indicates a substantial degradation in performance when the feature is permuted molnar2025interpretable. However, the notion of “large” depends on the scale of the evaluation metric. Because the raw magnitudes of $\overline{\text{PFI}}_{j}$ depend on the dataset and model, they are not directly comparable across problems. To provide a meaningful scale, we normalize by the importance of a random noise feature, $\overline{\text{PFI}}_{\text{noise}}$, to provide a noise-normalized $\widetilde{\text{PFI}}_{j}$ in eq 15.
@@
A $\widetilde{\text{PFI}}_{j}\approx 1$ implies that feature $j$ is no more predictive than random noise, while $\widetilde{\text{PFI}}_{j}>1$ indicates predictive power beyond noise. For example, $\widetilde{\text{PFI}}_{j}=5$ means that feature $j$ is five times as informative as the noise baseline. As shown in Figure 8, the anisotropy-based descriptors IAMS and QAMS dominate the model’s predictive performance, consistent with their direct link to polarization-dependent spectral asymmetries discussed in Section 2. The modest importance of geometric and chemical descriptors such as the normalized space group, density, and DAMS indicates that the model captures secondary contributions from the structural environment. The slight advantage of IAMS over QAMS likely reflects reduced uncertainty in quantities derived directly from computed spectra compared to oxidation-state–dependent charge estimates.
3.3 Model Performance for 3dd Transition Metals
Figure 9: Per-element performance of the random forest regressor evaluated on the test set. Ordered from largest (top) to smallest (bottom) atomic number. (a) shows the coefficient of determination $R^{2}$ computed separately for each metal using the combined (unbalanced) training dataset, and (b) shows the corresponding $R^{2}$ values obtained using the balanced dataset. In both panels, the model is trained on all elements and evaluated on element-specific subsets of the test data.
To better understand the limitations of the model, we analyzed its performance on a per-element basis. Specifically, we computed the coefficient of determination, $R^{2}$, for each individual metal using subsets of the test data corresponding to a single element, while keeping the random forest regressor (RFR) trained on the full dataset. The resulting per-element $R^{2}$ values are shown in Figure 9a. The coefficient of determination is defined in eq 16,
@@ 3.3 Model Performance for 3dd Transition Metals
Among all elements, Sc exhibits the poorest performance, with an $R^{2}$ value close to zero, indicating that the model provides little improvement over a naive mean predictor for this element. Figure 10 shows predicted versus true SAMS values for the Sc Figure 10a and Cu Figure 10b subsets of the test data. Showcasing the difference between a subset of the data that the model performs well on and one that it does not.
For Sc Figure 10a, the coefficient of determination is $R^{2}\approx 0.01$, which corresponds to a ratio of squared errors $\approx 0.99$. This indicates that the model reduces the error relative to a mean predictor by only about 1%. The small but positive $R^{2}$ value suggests that the model predictions are slightly shifted relative to the Sc-specific mean, consistent with a bias toward the global mean of the full multi-element dataset. In contrast to Sc, Cu Figure 10b exhibits excellent performance, with an $R^{2}$ of $0.851$, demonstrating that the model can accurately learn patterns when sufficient representative data is present.
Next, we investigate if the poor performance of Sc may be due to under representation of this element in the Materials Project database. To address this, we constructed a balanced data set with an equal number of compounds (444 compounds from each element, see Figure 6a). The training and testing sets were also split such that there was an even distribution of each element, with 354 samples in the training, and 90 samples in the test set. The results of this are presented in Figure 9b, where the performance for Sc significantly worsens, suggesting that its poor predictive behavior likely stems from intrinsic differences in the characteristics of Sc compared to the other 3dd transition metals rather than data imbalance.
@@ 4 Conclusion
In this work we demonstrated that a random forest regressor model is able to accurately predict a continuous anisotropy parameter using structural and chemical information extracted from crystallographic structures. This result provides an important step in creating new opportunities for machine learning driven materials discovery, as well as the design of materials with targeted spectroscopic properties. The model generalizes well to any crystalline 3dd transition metal structure (with the exception of Sc), but could easily be adapted to treat molecular systems as well by replacing or removing the normed space group feature.
@@
Future work includes extending our machine learning model to predict anisotropy in other x-ray methods, such as XAS. In particular, many of the structural and chemical descriptors that drive anisotropy in XES are also expected to influence absorption edge features, suggesting that the framework developed here could be transferable across spectroscopy techniques. It may be possible to apply a transfer learning approach so that the already trained XES anisotropy model can be fine-tuned for predicting anisotropy in a different technique, which could dramatically reduce the computational cost compared to training a new model from scratch.